OpenStudy (anonymous):

In a poll, respondents were asked whether they had ever been in a car accident. 157 respondents indicated that they had been in a car accident and 117 respondents said that they had not been in a car accident. If one of these respondents is randomly selected, what is the probability of getting someone who has been in a car accident?

OpenStudy (kropot72):

The total number of respondents = 157 + 117 = 274 \[P(InAccident)=\frac{157}{274}\]

OpenStudy (anonymous):

Translates to about 57.3%